### v3

# Time Entangled Quantum Blockchain with Phase Encoding for Classical Data

Abstract
With rapid advancements in quantum computing, it is widely believed that there will be quantum hardware capable of compromising classical cryptography and hence, the internet and the current information security infrastructure in the coming decade. This is mainly due to the operational realizations of quantum algorithms such as Grover and Shor, to which the current classical encryption protocols are vulnerable. Blockchains, i.e., blockchain data structures and their data, rely heavily on classical cryptography. One approach to secure blockchain is to attempt to achieve information theoretical security by defining blockchain on quantum technologies. There have been two conceptualizations of blockchains on quantum registers: the time-entangled Greenberger-Horne-Zeilinger (GHZ) state blockchain and the quantum hypergraph blockchain. On our part, an attempt is made to conceptualize a new quantum blockchain combining features of both these schemes to achieve the absolute security of the time-temporal GHZ blockchain and the scalability and efficiency of the quantum hypergraph blockchain in the proposed quantum blockchain protocol.
Keywords Quantum key distribution $\cdot$ Quantum networks $\cdot$ Quantum communication $\cdot$ Quantum cryptography $\cdot$ Quantum channels $\cdot$ Quantum entanglement $\cdot$ Quantum state $\cdot$ Blockchains $\cdot$ Consensus protocol $\cdot$ Authentication $\cdot$ Digital signatures $\cdot$ Cryptography $\cdot$ Security $\cdot$ Data integrity $\cdot$ Scalability $\cdot$ Quantum blockchain $\cdot$ Quantum hypergraph $\cdot$ Temporal entanglement $\cdot$ GHZ state $\cdot$ Information theoretical security
1 Introduction
  • Absolute Security / Unconditional Security: Security that cannot be compromised using unlimited computational capabilities. Classical security is conditional on the limitations of computational capabilities: dependent on the hardness of a computationally intractable mathematical problem. Information-theoretic security is the idea that needs to be propagated.
  • Temporal GHZ State: A GHZ state where particles are entangled not only in space but also in time.
  • Phase Encoding: Mapping the classical block into a phase angle and adding that phase to Bell and GHZ states.
  • Efficiency and Scalability:: the time-entangled GHZ blockchain encodes a classical block of 2 bits to a quantum block of 2 qubits. Ours is more efficient in the sense that a classical block of arbitrary size in encoded to a quantum block of 2 qubits. Therefore, due to this efficiency, we consider this protocol more scalable since it yields far less quantum information than the original GHZ blockchain [1] for a given amount of classical data. However, not as efficient as the quantum hypergraph blockchain [2], which encodes a classical block of arbitrary size to a quantum block of 1 qubit.
1.1 Classical Blockchain
A blockchain is an immutable ledger where data is stored on data structures known as blocks connected by hash pointers generated in chronological order. It is essentially a trustless peer-to-peer network [3], where time-stamped information (transactions in a cryptocurrency) is drawn from a pool of transactions and stored as a merkle tree in a block by a block creator: one of the nodes of the network chosen via a consensus protocol. These data are time-stamped (and encrypted if necessary) records: data about the past. It is a decentralized network that facilitates transactions/data without the necessity of a central party to authorize/authenticate its data and processes. Validated timestamped records that were put into the network in a given period by network nodes are packaged into a block and appended to the previous period’s block via a hash function [4]. The new block stores the hash of the previous block. The most famous use case of blockchain is Bitcoin [3]. A blockchain stores data on which all nodes have reached a consensus (agreement). There are many such consensus protocols, among which the most popular is PoW (Proof of Work). A block creator chosen by consensus validates the transactions he has put into the block and sends the block to the network where nodes validate the block and its transactions and append it to their local copy of the blockchain. This technology has found usage in various fields such as medicine [5] and social media [6]. The decentralized nature and the links formed by hash functions between blocks make it computationally impossible for an adversary to mutate a past record; if a block is mutated, its hash has to be recalculated and changed in the following block, recursively having to recalculate all the hashes for blocks following the mutated block, and since all nodes have a copy of the chain, the adoption of the mutated chain must be ensured as well. This requires the attacker to control more than half of the hash rate of the network, thus creating the longest local chain. The longest chain is eventually adopted by the network. Therefore, tampering with a block, due to the high degree of confidence of hash functions, effectively invalidates the blocks after the block in question. Therefore, the older the timestamp on a record in a blockchain, the safer it is: [7]
1.2 Quantum Blockchain
With the advent of quantum computing, algorithms such as Shor [8] and Grover [9] have shown themselves to be capable of compromising classical cryptography and hence, classical blockchain. There are applications and extensive research on quantum-immune classical blockchains based on quantum-immune classical cryptography. [10]. However, some of these schemes are broken by classical approaches and continuous research happens on new quantum algorithms to break these schemes as well. Therefore a sensible approach is to use quantum technologies to implement technical aspects of blockchain, the pinnacle of which is to define and store blockchains on a quantum register. These are known as quantum blockchains. An extensive analysis of quantum blockchains is in [7]. There are two main conceptualizations of quantum blockchains. The first is Del Rajan and Matt Visser’s [1] quantum blockchain on the temporal GHZ (Greenberger-Horne-Zeilinger) state. The other is Shreya Banerjee’s [2] blockchain on a quantum hypergraph. The GHZ blockchain is absolutely (information-theoretic security) secure in its temporal entanglement, while the hypergraph blockchain is scalable in the sense that it is efficient since a quantum block is equivalent to the data stored in a usual classical block. We have attempted to conceptualize a new blockchain with both these properties combining features of both these conceptualizations. The structure of this work is as follows. The GHZ blockchain and the quantum hypergraph blockchain are discussed. Then the security details of the GHZ blockchain and the efficiency of the hypergraph blockchain are outlined. The new quantum blockchain protocol is described: the blockchain structure, quantum network, security, consensus, and authentication. The chapter beyond that discusses future work to be done and comes before the concluding chapter.
2 Time Temporal GHZ State Blockchain
This blockchain was conceptualized in [1] by Del Rajan and Matt Visser. Entanglement is a phenomenon in which a compound state cannot be presented as a tensor of the states of the particles involved. It is the non-classical correlation between spatially distant particles that was referred to as "spooky action at a distance" [11] by Einstein. It is the foundation of quantum communication such as teleportation [12] and superdense coding. Progress toward a quantum internet [13] is being made on these technologies and quantum phenomena. An entangled quantum state is a compound state that cannot be given as a tensor of the states of quantum particles involved. A bipartite entangled state $|\psi_{ab}\rangle$ of particles $a$ and $b$ conforms to the following
$|\psi_{ab}\rangle\neq|a\rangle\otimes|b\rangle$
where $|a\rangle$ and $|b\rangle$ are individual qubit states.
multipartite GHZ (Greenberger–Horne–Zeilinger)states[14, 15] are states in which all subsystems (particles) contribute to the shared entangled property. This is used to conceptualize a chain. A concept from superdense coding where a classical 2-bit string $xy$ is encoded into a bell state by the following.
[…]
In this scheme, a classical block is two bits. The encoding procedure converts each classical block $r_{1}r_{2}$, into a temporal Bell state [16], generated at a particular time ($t=0$)
$|B_{r_{1}r_{2}}\rangle^{0,\tau}=\frac{1}{\sqrt{2}}(|0^{0}\rangle|r_{2}^{\tau}\rangle+(-1)^{r_{1}}|1^{0}\rangle|\tilde{r}_{2}^{\tau}\rangle)$
The superscript in the kets is the time at which the photon is absorbed; this provides a timestamp for the block. The authors of [16] experimentally generated such temporal Bell states. They presented spatial bell states with polarized photons where $v_{a}(h_{a})$ represent the vertical (horizontal) polarization in spatial mode $a(b)$. To create the temporally entangled states, consecutive pairs of spatially entangled pairs are generated at defined intervals separated by interval $\tau$.
$|\psi^{-}\rangle_{a,b}^{0,0}\otimes|\psi^{-}\rangle_{a,b}^{\tau,\tau}=(|h^{0}_{a}v^{0}_{b}\rangle-|v^{0}_{a}h^{0}_{b}\rangle)\otimes(|h^{\tau}_{a}v^{\tau}_{b}\rangle-|v^{\tau}_{a}h^{\tau}_{b}\rangle)$
A delay line was added to one photon of each pair in [16].
$|\psi^{-}\rangle_{a,b}^{0,\tau}|\psi^{-}\rangle_{a,b}^{\tau,2\tau}=\frac{1}{2}\Big(|\psi^{+}\rangle_{a,b}^{0,2\tau}|\psi^{+}\rangle_{a,b}^{\tau,\tau}-|\psi^{-}\rangle_{a,b}^{0,2\tau}|\psi^{-}\rangle_{a,b}^{\tau,\tau}$-
$|\phi^{+}\rangle_{a,b}^{0,2\tau}|\phi^{+}\rangle_{a,b}^{\tau,\tau}+|\phi^{-}\rangle_{a,b}^{0,2\tau}|\phi^{-}\rangle_{a,b}^{\tau,\tau}\Big).$
The photons absorbed at $t=0$ and $t=2\tau$ are entangled when Bell projection is carried out on two photons at $t=\tau$. The former pair of photons never even coexists. In the design of the blockchain in [1], 2-bit blocks are encoded into temporal Bell states, where photons are absorbed at respective times of encoding. The first three blocks,
$|\beta_{00}\rangle^{0,\tau},\quad|\beta_{10}\rangle^{\tau,2\tau},\quad|\beta_{11}\rangle^{2\tau,3\tau}$
These states (blocks) are chained together through entanglement in time. This is done by recursively projecting these Bell states into a growing GHZ state an entanglement source, a delay line, and a polarizing beam splitter (PBS). The following is an example of two Bell pairs forming a four qubit GHZ state
$|\psi^{+}\rangle_{a,b}^{0,0}\otimes|\psi^{+}\rangle_{a,b}^{\tau,\tau}\overset{delay}{\longrightarrow}|\psi^{+}\rangle_{a,b}^{0,\tau}\otimes|\psi^{+}\rangle_{a,b}^{\tau,2\tau}=\frac{1}{2}(|h_{a}^{0}v_{b}^{\tau}\rangle+$
$|v_{a}^{0}h_{b}^{\tau}\rangle)\otimes(|h_{a}^{\tau}v_{b}^{2\tau}\rangle+|v_{a}^{\tau}h_{b}^{2\tau}\rangle)$
$\overset{PBS}{\longrightarrow}\frac{1}{2}\left(|h_{a}^{0}v_{b}^{\tau}h_{a}^{\tau}h_{b}^{2\tau}\rangle+|v_{a}^{0}h_{b}^{\tau}h_{a}^{\tau}v_{b}^{2\tau}\rangle\right)=|GHZ\rangle_{0,\tau,2\tau}$
According to [1], "Entanglement exists between the four photons that propagate in different spatial modes and exist at different times." From $t=0$, the state of the blockchain at $t=n\tau$:
$|GHZ_{r_{1}r_{2}...r_{2n}}^{0,\tau,\tau,2\tau,2\tau..,(n-1)\tau,(n-1)\tau,n\tau}\rangle=\frac{1}{\sqrt{2}}(|0^{0}r_{2}^{\tau}r_{3}^{\tau}...r_{2n}^{n\tau}\rangle+$
$(-1)^{r_{1}}|1^{0}\bar{r}_{2}^{\tau}\bar{r}_{3}^{\tau}...\bar{r}_{2n}^{n\tau}\rangle)$
Subscripts on the left side denote the concatenated string of all the blocks, i.e all the data stored on the chain. At $t=n\tau$, only one photon remains. The following is to quote [1].
It is important to note that at this stage of development, we are advocating a conceptual mathematical design for a new quantum information technology. It should be viewed as analogous to early quantum algorithms (Deutsch’s algorithm [17], Deutsch-Jozsa algorithms [18]). In the 1980s, the engineering considerations for quantum computers were not taken into account. In the 1990s, when Shor’s algorithm and Grover’s algorithm were developed, the experimental realization of quantum computers was almost seen as an impossible project. Nonetheless, their work was certainly of interest to the wider community [19].
3 Quantum Blockchain on a Weighted Hypergraph
This scheme was presented in [2] by Shreya Banerjee, A. Mukherjee, and P. K. Panigrahi. Quantum hypergraphs [20] are a group of highly entangled multiparty quantum states. Vertices are quantum particles where edges represent entanglement. Edges encompass more than two vertices and are known as "hypergraphs." With a quantum hypergraph with $k$-hyperedge (a hyperedge connecting $k$ qubits) and n1n-1 vertices, a corresponding quantum state can be prepared [20]. To add the $n$th qubit to the state, initialize the qubit to the Hadamard state: $(|0\rangle+|1\rangle)/\sqrt{2}$. A controlled $Z$ operation is performed with the existing n1n-1 qubits as control and $n$ as the target. For a mathematical hypergraph with five vertices, A, B, C, D, E, a 3-hyperedge over vertices A, B, and C; and a 5-hyperedge over all five vertices, the state is expressed as following.
$|\psi\rangle=C^{2}_{(A,B,C,D,E)}ZC^{2}_{(A,B,C)}Z|+\rangle^{\otimes 5}$
The entanglement of a hypergraph state and its properties have been discussed in [21, 20, 22].
[21] describes how local unitary operations carried through classical communication (LOCC) does not alter the entanglement of the state under consideration. It is shown in [21] that local applications of unitary Pauli operations on the $n$th qubit one can remove all the (N1)(N-1) edges for the special case where an n-hypergraph ($n$-qubit) contains only an $n$ hyperedge. As presented in [22, 23], a weighted hypergraph is when hyperedge carries weights. [22, 24] introduced weighted quantum hypergraphs as locally maximally entangleable (LME) states. In the notation below, $|x\rangle$ is the computational basis, and $f(x)$ is a real number.
$|\psi\rangle=\frac{1}{\sqrt{2^{N}}}\sum_{x\in(0,1)^{N}}e^{i\pi f(x)}|x\rangle.$
In the proposed blockchain, the classical ledger and cryptographic functions have been replaced by the weighted quantum hypergraph and entanglement. They have used the phases (weights) on the hyperedges to encode classical information. Each classical block is a string $P_{i}$ (iith classical block) of bits. The block creator initializes the qubit (one qubit represents one quantum block) to $(|0\rangle+|1\rangle)/\sqrt{2}$. Then via a secret bijective function chosen known only by him, $P_{i}$ is converted into the phase $\theta_{P_{i}}$, and the following rotation is applied to the qubit to obtain the $i$th quantum block.
$|\psi_{i}\rangle=S(p)|\psi\rangle=\frac{1}{\sqrt{2}}\begin{bmatrix}1&0\\ 0&e^{i\theta_{p}}\end{bmatrix}|\psi\rangle=\frac{|0\rangle+e^{i\theta_{p}}|1\rangle}{\sqrt{2}},$
$\theta_{P_{i}}\in(0,\pi/2)$ is $f_{i}(P)$ the output of the aforementioned bijective function $f_{i}$. $\theta$ should satisfy $\sum_{i}^{\infty}p_{i}<\pi/2$. The qubit $|\psi_{i}\rangle$ (quantum block i) now carries the information in classical n block $P_{i}$. Conditions on the phase are of critical importance as they ensure the entanglement of the hypergraph. These conditions are part of the consensus. Constraints include $\theta_{P_{i}}=\theta_{P_{1}}/2^{(i-1)}$. where $\theta_{P_{1}}$ is the phase of the first block. The infinite sum of all the phases
$\sum_{i=1}^{\infty}\theta_{p_{i}}=\sum_{i=1}^{\infty}\frac{1}{2^{i-1}}\theta_{p_{1}}.$
converges to $2\theta_{P_{1}}$. Therefore, to ensure $\sum_{i}^{\infty}p_{i}<\frac{\pi}{2}$, $\theta_{P_{1}}$ needs to be less than $\frac{\pi}{4}$.
There can be many such series.
$\sum_{i=1}^{\infty}\theta_{p_{i}}=\sum_{i=1}^{\infty}\frac{1}{n^{i-1}}\theta_{p_{1}}.$
where $n\in\mathbb{N}\setminus\{1\}$. The series converges to $\frac{n}{n-1}$. Consensus can therefore be defined with any such series with an appropriate $\theta_{P_{1}}$. After the consensus execution, peers add the $n$th block to their local chain of n1n-1 blocks. A $C^{(n-1)}Z$ is applied as before with the new block as the target to create the new $n$-hyperedge and a local Pauli-$X$ on the new block to remove the previous n1n-1-hyperedge.
4 New Quantum Blockchain
4.1 Security and Scalability
The GHZ blockchain [1] has unconditional security in the sense that its entanglement is temporal in time. In just the spatial GHZ case, tampering (measuring) with any photon would result in the full local copy of the blockchain being invalidated (collapse) immediately. The temporal GHZ blockchain adds much greater security in that the attacker cannot even attempt to access the previous blocks since they no longer exist. The adversary could try to tamper with the last remaining photon, which would invalidate the full state. Hence the absolute security of entanglement in time. The scalability of these temporal GHZ states was considered in [16]. They have stated " any number of photons are generated with the same setup, solving the scalability problem caused by the previous need for extra resources. Consequently, entangled photon states of larger numbers than before are practically realizable." In any case, the proposed blockchain encodes only one classical bit per qubit. However, the hypergraph blockchain [2] has encoded an entire classical block into a single qubit, showing theoretical promise in scalability: scalable due to efficiency from encoding an entire classical block into a single qubit. An attempt is made by us to combine this absolute security from the time-temporal GHZ blockchain and the efficiency from the hypergraph blockchain’s classical data encoding.
4.2 Blockchain Protocol
In the new blockchain, for the classical block i, i.e. $P_{i}$, phase $\theta_{P_{i}}$ is calculated by the block creates much the same as in the hypergraph blockchain via a secret bijective function known only to the block creator. However, the outcome of this function is this phase ++ a classical $2$-bit string $r_{1}r_{2}$. There are two qubits per block, initially in the ground state $|00\rangle$. The following state is then obtained, i.e., the corresponding Bell state for the classical bit string.
$|\psi\rangle=\frac{|0r_{2}\rangle+(-1)^{r_{1}}|1\bar{r_{2}}\rangle}{\sqrt{2}}$
Apply the following rotation to obtain the block if $r_{1}r_{2}=00$ or $r_{1}r_{2}=10$.
$|\psi_{i}^{\prime}\rangle=S(p)|\psi\rangle=\frac{1}{\sqrt{2}}\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&e^{i\theta_{p}}\end{bmatrix}|\psi\rangle=$
$\frac{|0r_{2}\rangle+e^{i\theta_{p_{i}}}(-1)^{r_{1}}|1\bar{r_{2}}\rangle}{\sqrt{2}}$
Otherwise, apply the following rotation.
$|\psi_{i}^{\prime}\rangle=S(p)|\psi\rangle=\frac{1}{\sqrt{2}}\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&e^{i\theta_{p}}&0\\ 0&0&0&1\end{bmatrix}|\psi\rangle=$
$\frac{|0r_{2}\rangle+e^{i\theta_{p_{i}}}(-1)^{r_{1}}|1\bar{r_{2}}\rangle}{\sqrt{2}}$
Then the iithe block is
$|\psi_{i}\rangle=\frac{|0r_{2}\rangle+e^{i\theta_{p_{i}}}(-1)^{r_{1}}|1\bar{r_{2}}\rangle}{\sqrt{2}}$
In the temporal case, the following state is the block of the new scheme. [25] suggests that temporal Bell states and GHZ states can have a phase, but they emphasize that there is no need for sensitive phase accuracy: only an overlap of pulse envelopes is required for interference. Strict phase control is not necessary or accounted for in their framework, but we have assumes strict phase control.
$|\psi_{i}\rangle^{0,\tau}=\frac{|0r_{2}\rangle^{0,\tau}+e^{i\theta_{p_{i}}}(-1)^{r_{1}}|1\bar{r_{2}}\rangle^{0,\tau}}{\sqrt{2}}$
Suppose we have the first three blocks following this notation.
$|\phi_{1}\rangle^{0,\tau}=\frac{|0r_{1_{2}}\rangle^{0,\tau}+e^{i\theta_{p_{1}}}(-1)^{r_{1_{1}}}|1\bar{r_{1_{2}}}\rangle^{0,\tau}}{\sqrt{2}}$
The bijective function for the genesis block is chosen such that $r_{1_{1}}r_{1_{2}}$ is either $00$ or $01$ to keep the phase of the blockchain below $\pi/2$.
$|\phi_{2}\rangle^{\tau,2\tau}=\frac{|0r_{2_{2}}\rangle^{\tau,2\tau}+e^{i\theta_{p_{2}}}(-1)^{r_{2_{1}}}|1\bar{r_{2_{2}}}\rangle^{\tau,2\tau}}{\sqrt{2}}$
$|\phi_{3}\rangle^{2\tau,3\tau}=\frac{|0r_{3_{2}}\rangle^{2\tau,3\tau}+e^{i\theta_{p_{3}}}(-1)^{r_{3_{1}}}|1\bar{r_{3_{2}}}\rangle^{2\tau,3\tau}}{\sqrt{2}}$
These will be fused together using the fusion process in [25], i.e. projected recursively onto a growing time temporal GHZ state which is the blockchain using a PBS and a delay line in addition to the entangled pair generation equipment. They have suggested that strict phase control is not necessary but we have assumed it since precise phase is an integral part of our scheme. [26] also implies phase in temporal GHZ states. The presence of a well-defined phase in these states depends on whether coherence can be established between different temporal modes, which we have assumed. We have assumed temporal coherence and synchronization of the time modes [27], precise control of the phase evolution of each temporal mode, minimal noise and decoherence to maintain phase stability[28], perfect entanglement [29] and coherence between the temporal modes, external reference phase to define the phase, etc.
$|chain\rangle^{0,\tau,\tau,2\tau,2\tau,3\tau}_{r_{1_{1}}r_{1_{2}}r_{2_{1}}r_{2_{2}}r_{3_{1}}r_{3_{2}}}=\frac{1}{\sqrt{2}}(|0^{0}r_{1_{2}}^{\tau}r_{2_{1}}^{\tau}r_{2_{2}}^{2\tau}r_{3_{1}}^{2\tau}r_{3_{2}}^{3\tau}\rangle+$
$(-1)^{r_{1_{1}}}e^{i(\theta_{P1}+\theta_{P2}+\theta_{P3})}|1^{0}\bar{r}_{1_{2}}^{\tau}\bar{r}_{2_{1}}^{\tau}\bar{r}_{2_{2}}^{2\tau}\bar{r}_{3_{1}}^{2\tau}\bar{r}_{3_{2}}^{3\tau}\rangle)$
The general state of the blockchain at $t=n\tau$
$|chain\rangle^{0,\tau,\tau,2\tau,\dots(n-2)\tau,(n-1)\tau,(n-1)\tau,n\tau}_{r_{1_{1}}r_{1_{2}}r_{2_{1}}r_{2_{2}}\dots r_{n-1_{1}}r_{n-1_{2}}r_{n_{1}}r_{n_{2}}}=$
$\frac{1}{\sqrt{2}}(|0^{0}r_{1_{2}}^{\tau}r_{2_{1}}^{\tau}r_{2_{2}}^{2\tau}\dots\bar{r}_{n-1_{1}}^{(n-2)\tau}\bar{r}_{n-1_{2}}^{(n-1)\tau}\bar{r}_{n_{1}}^{(n-1)\tau}\bar{r}_{n_{2}}^{n\tau}\rangle+$
$(-1)^{r_{1_{1}}}e^{i(\theta_{P1}+\theta_{P2}+\dots\theta_{P(n-1)}+\theta_{Pn})}|1^{0}\bar{r}_{1_{2}}^{\tau}\bar{r}_{2_{1}}^{\tau}\bar{r}_{2_{2}}^{2\tau}\dots$
$\bar{r}_{n-1_{1}}^{(n-2)\tau}\bar{r}_{n-1_{2}}^{(n-1)\tau}\bar{r}_{n_{1}}^{(n-1)\tau}\bar{r}_{n_{2}}^{n\tau}\rangle)$
where $r_{1_{1}}r_{1_{2}}$ is either $00$ or $01$.
There are constraints on the phase similar to the hypergraph blockchain as part of the consensus. Consensus will be discussed later. The first is $\sum_{i=1}^{\infty}\theta_{P_{i}}<\frac{\pi}{2}$, following [2], in order to create a basis for consensus.
$\theta_{P_{i}}=\ \frac{\theta_{P_{1}}}{n^{(i-1)}}$
where $n\in\mathbb{N}\setminus\{1\}$. $\theta_{P_{1}}$ is shared with the entire network by the creator of the genesis block along with $r_{1_{1}}r_{1_{2}}$. $\sum_{i=1}^{\infty}\theta_{P_{i}}=\sum_{m=1}^{\infty}\frac{\theta_{P_{1}}}{(n-1)^{m}}=\frac{n}{n-1}\theta_{P_{1}}$. The upper bound is $\pi/2$. Therefore when $n=2$, it needs to be $\theta_{P_{1}}<\pi/4$. A suitable $\theta_{1}$ needs to be chosen with the chosen $n$.
Similar to the GHZ blockchain [1] and the hypergraph blockchain [2], it must be understood that the blockchain scheme proposed here is purely and entirely conceptual, and practical implications have not been accounted for in the least. This work must be viewed as a root for extensive future research on the refinement, extension, and correction of this theoretical model or the practical realization of this blockchain scheme. When algorithms Shor [8] and Grover [19] were conceptualized, they were entirely theoretical for quantum computers were far from realization; this blockchain must be viewed from the same perspective.
4.3 Quantum Network
Similar to both [2] and [1], the network assumes both a classical network and a quantum network [30], where each pair of nodes has an unsecured classical channel and a secure quantum channel. The distribution of quantum blocks requires quantum communication channels. A quantum key distribution method of choice can be used to secure classical communication. We have assumed this quantum network with a QKD (Quantum Key Distribution)layer, much like the paper that introduced the $\theta$-protocol, where they have made the same exact assumption. The quantum communication (quantum state, i.e., public keys, distribution via quantum communication channels) layer will be utilized for the signature protocol from Gottesman [31]. This quantum public key signature scheme is useful for signing transactions and, more importantly, for authenticating classical communication. In addition to the utilization by this signature, the quantum channels will be used to perform QKD for the symmetric encryption of all classical communication necessary in consensus; transaction, signature, and the other data transmittance across the network; etc. The trusted key distribution authority they have suggested or any other method of quantum state distribution can be used for quantum public key distribution. Quantum network infrastructure to facilitate operations of the signature, such as swap tests, key set extensions, global pure state [31], and quantum public key storage is assumed in this regard. However, a centralized [31] key distribution may not be ideal since blockchain is decentralized.
Space-time and space effects in satellite-based quantum networks [32] is a topic in research along with quantum network [33, 34, 35] advancement in general.
4.4 Consensus
Consensus, at a high level, is similar to the hypergraph blockchain. To validate the phase, the block creator distributes a large number of copies of the state of the $m$th quantum block to the entire network, i.e. the state $\frac{|0r_{m_{1}}\rangle+(-1)^{r_{m_{1}}}e^{i\theta_{m}}|1\bar{r}_{m_{1}}\rangle}{\sqrt{2}}$, is distributed to each network node without violating the no-cloning theorem, and each node can use all except one to check the validity and the other to append the next block to their local chain. The classical string $r_{m_{1}}r_{m_{2}}$ (to be securely stored) is distributed by classical channels secured via QKD. This state distribution is done via the mutual quantum channels. At this stage of the design, following [1], we assumed that newly generated blocks are spatial GHZ states since entangling these qubits in time at this stage of the design process is unnecessary and is left for future work. The block producer is chosen randomly, utilizing a low-level algorithm based on Quantum Random Number Generation (QRNG) [36].
$|\phi^{\prime}_{m}\rangle=\frac{|0r_{m_{1}}\rangle+(-1)^{r_{m_{1}}}e^{i\theta_{m}}|1\bar{r}_{m_{1}}\rangle}{\sqrt{2}}$
[…]
In classical blockchain, there are genres based on the degree of centralization of governance and access [37]: private (permissioned), public (open), based in terms of access and consortium, centralized, and fully centralized. There are different consensus protocols [38] for these different types, mainly being of two categories: proof-based, and vote-based. Not all classical, post-quantum, and quantum blockchain frameworks satisfy Byzantine fault tolerance. Hyperledger Fabric [39] early versions and Nakamoto’s [3] PoW are not Byzantine fault tolerant: [40]. With that knowledge, we believe it is not sensible to expect a full security proof for byzantine fault tolerance for our blockchain, which is still in early conceptualization.
###What needs to be emphasized is that all aspects of this blockchain framework need extensive low level and high-level research/development to be realized practically. This includes the blockchain protocol, consensus, security, authentication, etc.
The temporal GHZ state in this blockchain framework involves an entanglement between photons that do not share simultaneous coexistence [16], yet they share nonclassical measurement correlations. It can be interpreted as a way to link records in the current block to records not $of$ the past but $in$ the past.
4.5 Security and Integrity
All nodes share the measurement outcomes, judgement on the block creator’s honesty, and $r_{i_{1}}r_{i_{2}}$ they received from the block creator with all other nodes. Then based on other nodes’ judgements and information, a node shares their final verdict on the admissibility of the block. These validating nodes can compare the classical 2-bit strings $r_{1}r_{2}$ shared by other nodes and determine if there are signs of the block creator having distributed different bit strings to different nodes. If such a receiving node is dishonest, they could either report the block creator as dishonest when the new block is valid or report honest when the new block’s phase is invalid. In any case, as long as the majority, i.e. more than a half is honest, their judgment on the admissibility of the block will be accepted; this is due to the fact that all honest nodes will come to the same conclusion about the admissibility. All nodes with a different judgment would be deemed untrustworthy.
In the quantum blockchain, the traditional time-stamped blocks and hash functions connecting them are replaced by a temporal GHZ state with a growing phase, utilizing entanglement over time. Hence, the sensitivity to tampering is greatly amplified. If a single block/qubit is measured, the entire local copy of the blockchain is compromised due to the entanglement, whereas, in a classical blockchain, only the blocks following the tampered one are affected (hash pointers), leaving the system susceptible to vulnerabilities. In classical blockchains, it is often claimed that the further back a block is in the chain, the more "secure" it becomes. This is because tampering with a block earlier in the chain invalidates more data, thus strengthening security.
4.5.1 Theoretical Tamper Sensitivity/Security of the GHZ Blockchain in Comparison to the Quantum Hypergraph Blockchain[2]
It is important to note that the weighted hypergraph states used in [2] do not possess the same degree of multipartite entanglement as maximally entangled GHZ-type states. A phased GHZ state,
$|{\rm GHZ}_{\phi}\rangle=\frac{1}{\sqrt{2}}\left(|0\ldots 0\rangle+e^{i\phi}|1\ldots 1\rangle\right),$
remains maximally entangled for any value of the phase $\phi$, and is extremely fragile under local measurements: a projective measurement on any single qubit collapses the entire state and destroys the entire multipartite entangled state.
## Start of [41]
The "Non"-Maximal Entanglement of Weighted Hypergraph States[41]
According to [41], a weighted graph state (WGS), a subclass of Weighted Quantum Hypergraph States (WQHS), is generated when controlled-phase (CP) gates with an angle $\phi\neq\pi$ are used, instead of the maximally entangling Controlled-Z (CZ) gates ($\phi=\pi$). This implies that the resulting state is generally not maximally entangled. This paper, "Measurement-based quantum computation on weighted graph states with arbitrarily small weight,"[41] explicitly addresses the nature of the entangling gates and the resulting state:
"Weighted graph states are a natural generalization of graph states, which are generated by applying controlled-phase gates, instead of controlled-Z gates, to a separable state. … One such experimental constraint is that only controlled-phase (CP) gates with angle $\phi$ are available, instead of controlled-Z (CZ) gates ($\phi=\pi$). Then, a generated state becomes a so-called **weighted graph state** (WGS), instead of a graph state." [41]
[41] Continuation
They further classify the entanglement inherent to these states, stating that WGS are the first example of a universal resource created this way:
"To our knowledge, this is the first example of universal resources prepared with only **non-maximally entangling gates** and has potential applications to weakly interacting systems, such as photonic systems." [41]
This use of non-maximally entangling gate to prepare the WGS (and, more generally, WQHS with variable weights) suggests that the resulting states may not always be maximally entangled.
## End of [41]
In contrast to GHZ, the weighted hypergraph states employed in the hypergraph blockchain construction [2] have a significantly more distributed and non-uniform entanglement structure. As stated above, these states include amplitudes for all computational basis states,
$|\psi\rangle=\frac{1}{\sqrt{2^{N}}}\sum_{x\in\{0,1\}^{N}}e^{i\pi f(x)}|x\rangle,$
with phases determined by the pattern and weights of multiple $k$-hyperedges. Because the entanglement in such states is typically non-maximal and is spread across many local correlations, their response to local measurements is less catastrophic than that of a GHZ state. In general, measuring a single qubit of a weighted hypergraph state theoretically does not theoretically annihilate all global correlations, and the post-measurement state may still retain nontrivial entanglement among the remaining qubits.
Thus, from a theoretical standpoint, GHZ states are more fragile and more sensitive to tampering or measurement, whereas the weighted hypergraph states used in this protocol possess a resilient but non-maximal entanglement (less entangled) structure.
For instance, the 3-hyperedged, 3-qubit hypergraph(blockchain) from [2] is
$|\psi\rangle_{123}=X_{3}C_{12}^{2}C_{13}^{2}\,|\psi\rangle_{12}\otimes|\psi\rangle_{3}$
$\frac{1}{2\sqrt{2}}\Big(e^{i\theta_{p_{3}}}|000\rangle+|001\rangle+e^{i(\theta_{p_{2}}+\theta_{p_{3}})}|010\rangle+e^{i\theta_{p_{2}}}|011\rangle+e^{i(\theta_{p_{1}}+\theta_{p_{3}})}|100\rangle+e^{i\theta_{p_{1}}}|101\rangle+e^{i(\theta_{p_{1}}+\theta_{p_{2}}+\theta_{p_{3}})}|110\rangle-e^{i(\theta_{p_{1}}+\theta_{p_{2}})}|111\rangle\Big).$
, where it is intuitive that, for any $\theta_{p_{1}},\theta_{p_{2}},\theta_{p_{3}}$ chosen according to the consensus protocol in [2], this state is not as entangled as a GHZ state. It is worth emphasizing that although weighted hypergraph states belong to the class of LME states, this property does not in general imply maximal multipartite entanglement. In the sense used in this work, LME refers to the ability of a state to generate maximal entanglement between the system and a set of local ancillas via suitable local unitary operations. This is an operational notion that concerns the entangling power of the state under local control, rather than the intrinsic amount of global entanglement present in the multipartite system itself. In particular, the degree of genuine multipartite entanglement depends sensitively on the choice of hyperedge weights and the resulting phase structure. For generic weight assignments, the reduced density matrices of subsystems need not be maximally mixed, and standard entanglement measures need not attain their maximal values. Consequently, while the LME property ensures strong nonclassical correlations and local controllability, it does not guarantee that the corresponding weighted hypergraph states are maximally entangled in the multipartite sense.
Hence, a GHZ blockchain is theoretically more tamper/measurement-sensitive and secure than the hypergraph blockchain[2].
4.5.2 The Spatial GHZ Blockchain Security: Without the Temporal Aspect
The data is encoded into the phase of the GHZ state and the basis states themselves. Therefore, for an adversary to measure the phase, they will need to keep measuring the state in random basis as long as the length of the chain, $\theta_{P_{1}}$, classical 2-bit strings, or $n$ remain secrets. This randomness includes GHZ basis states computational basis states as well. Even if the chain is constructed by the adversary using those parameters (even if he obtains bit strings and the phases of blocks), without the knowledge of the bijective functions, it is not possible to obtain the data. Even if the data is obtained, i.e, the bijective functions are discovered, it is still impossible to mutate the data with that same bijective function intact, i.e, change the phase of the $i$th block - $\theta_{Pi}$ - and the classical string - $r_{i_{1}}r_{i_{2}}$ -, since this would result in a different phase than that which is predefined for the GHZ state and different basis states that don’t represent the classical bit strings $r_{i_{1}}r_{i_{2}}$. If the length of the chain is $m$, i.e. $2m$ qubits, the projection measurement basis to check validity will be
$\frac{1}{\sqrt{2}}(|0r_{1_{2}}r_{2_{1}}r_{2_{2}}\dots\bar{r}_{n-1_{1}}\bar{r}_{n-1_{2}}\bar{r}_{n_{1}}\bar{r}_{n_{2}}\rangle\pm$
$(-1)^{r_{1_{1}}}e^{i\theta_{expected}}|1\bar{r}_{1_{2}}\bar{r}_{2_{1}}\bar{r}_{2_{2}}\dots\bar{r}_{n-1_{1}}\bar{r}_{n-1_{2}}\bar{r}_{n_{1}}\bar{r}_{n_{2}}\rangle)$
with the rest of the basis set: the other $2^{2n}-2$ states that make an orthonormal basis (GHZ basis with an added phase) with these 2 states, where in the projection $\Omega,I-\Omega$}, $\Omega$ is the projection operator for the ++ state out of the two states above.
Also, in the above, $\theta_{expected}=\sum_{i=1}^{m}\theta_{m}=\sum_{i=1}^{m}\frac{\theta_{1}}{n^{(m-1)}}$, and $\theta_{P_{1}}$, i.e. $\theta_{1}$, is shared with the entire network by the creator of the genesis block. Since the phase is the phase of the whole compound GHZ state, when changing the phase of the $i$th block, it is suggested that $\theta_{Pi}$ is changed, and therefore the phase of the entire GHZ state is changed, not the phase of individual pairs of qubits or single qubits.
Measurement in the correct basis will reveal whether the phase and GHZ basis states are valid, i.e. the predefined phase and locally stored classical 2-bit strings. The outcome must be deterministic for the basis state the blockchain should be in with probability no less than $1$. Any other outcome reveals that the chain is invalid for the current length and the owner of that local copy can reconstruct the blockchain with the use of the length of the blockchain, $\theta_{P_{1}}$, 2-bit classical strings received from block creators, and $n$. If there is any suspicion of mutation, reconstruction can be done at any point. Depending on the exact application built on the blockchain, the secret bijective functions would need to be shared with the network; certain functions are disclosed to certain nodes.
In the temporal case, only the last remaining qubit, which holds the entire aggregated phase of the entire chain, can be modified in its phase or GHZ basis states: the above arguments and theory apply accordingly. For instance, the basis will account for the last qubit with the phase being the same as above.
With just a spatial GHZ state, measurement correlations are stronger than any classical blockchain. The chain being a maximally entangled state, any measurement on any of the qubits will lead to a collapse of the entire local copy. In this scenario, if an attacker alters any photon, the entire local copy of the blockchain is immediately invalidated (collapses), offering a clear advantage over classical systems, where only blocks after the tampered one are affected. The temporal GHZ blockchain offers even stronger security, as past photons no longer exist, preventing any attempt to access or tamper with them. An attacker can only alter the final remaining photon, which would immediately invalidate the entire state. This demonstrates that temporal entanglement provides a significantly greater security advantage than spatial entanglement.
A case by case analysis of security is still necessary and is left for future.
4.6 Local Unitary Operations on the blockchain
The spatially entangled blockchain is accounted for in this case. The idea of applying these transformation is to further solidify security by obscuring the actual state of the blockchain from attackers, making it impossible to mutate the chain, i.e., the phase, or extract data even with the knowledge of the initial phase, the length $m$ of the $2m$ qubit long chain, some and the bijective functions: the adversary wouldn’t be able to obtain the chain by reverse transformations or construct the chain and therefore the phase of the blockchain. Peers can apply local unitary operations to the subsystems of the quantum blockchain, which is a compound quantum state. In the case of a local unitary operation to a subsystem, the operation would always be reversible since unitary operations are reversible, but depending on the actual transformation applied, the phase, after the reversal, might not exactly be the same as the original phase. However, a global unitary operation such as a phase shift on the whole state would be reversible to the same initial blockchain. In any case, if the inverse transformation cannot restore the original state, with the knowledge of $n$, chain length $m$, all the $2$-bit strings, and $\theta_{P1}$, an honest peer would be able to reconstruct the blockchain.
In the temporal case, since only one qubit with the cumulative phase exists, any rotation or unitary transformation is possible and reversible.
5 Future Directions
[1] has discussed how their time temporal blockchain can be viewed as a way to influence the past in non-classical ways, i.e. a quantum networked time-machine. They believe it could lead to an information theoretical investigation into the nature of time: [42, 43]. We could state the same regarding our scheme, which is a temporal GHZ state with phase. Similarly to [1], "we speculate that a blockchain can be encoded into a different temporally entangled system, namely, the entanglement between the future and past in the quantum vacuum:[44, 45]. A realistic experimental proposal [46] suggests that it is possible to transfer this future past quantum correlation into qubits that do not simultaneously coexist, which is the resource needed for our current design."
6 Conclusion
We have described the two main quantum blockchain data structures: Temporal GHZ state blockchain [1] and quantum hypergraph blockchain [2]. The GHZ blockchain is absolutely secure in the sense that older blocks no longer exist [1] to be attacked, making it more secure from a theoretical standpoint than hypergraphs whose entanglement may not even be always as strongly correlated as the spatial case of the GHZ state. The temporal aspect then takes security to another sphere. The hyperghraph blockchain, with its phase encoding with bijective functions, where an entire classical block is encoded into a one-qubit quantum block, is much more efficient and therefore scalable than the GHZ state blockchain which has $2$ or $n$ qubit quantum blocks and encoding $m$-bit classical data takes $O(m)$ quantum bits. An attempt was made by us to formulate a new quantum blockchain structure with this absolute security in the temporal GHZ state and the efficiency and scalability in the hypergraph blockchain: a temporal GHZ blockchain with phase encoding. We have formulated a basic authentication and consensus model and a basic basic security analysis, all of which need thorough refining and research at the lower level even in the conceptual stage.
7 Limitations
It must be understood that this work is to merely convey a conceptual blockchain on a quantum register. Low-level, case-by-case attacker and security models, a quantum consensus algorithm tailored to the phased GHZ and Bell states in question must be constructed. This work is not intended to be a fully formalized, rigorous theoretical framework or an engineering-ready design of a blockchain.
  1. 1.
    Lack of full consensus protocol formalization. While we outline a high-level consensus mechanism, a fully detailed quantum consensus protocol with rigorous message flows, fault tolerance proofs, and an explicit adversarial model in a quantum network remains to be done. We invite enthusiastic academics to build upon our base. Similar to [1], the starting point for this development could consider the $\theta$-protocol [47].
  2. 2.
    Adversary and threat model constraints. The security analyses assume an idealized quantum network, limited error models, and simplified adversarial capabilities. A comprehensive, case-by-case attacker model (including quantum memory attacks, denial-of-service, network partitioning, and measurement coercion) is not fully treated.
  3. 3.
    Technology and implementation readiness. The proposal leverages time entanglement, phase encoding, and phased GHZ states. No empirical prototype or detailed engineering study on current photonic hardware, error rates, or scalability in practical deployment is included. All features adopted from [1] have been proven in the empirical sense as claimed by [1]. Feasibility of large-scale quantum-register blockchain networks remains to be demonstrated. All features adopted from [2] are empirically proven through their simulations. It remains to be empirically proven whether phase can be added to spatial and temporal GHZ states.
  4. 4.
    Scalability and resource analysis. While combining time-entanglement with phase encoding aims to enhance efficiency and scalability of the temporal GHZ blockchain [1], detailed quantitative resource estimates (e.g., qubit counts, entanglement generation overhead, decoherence thresholds, throughput/latency metrics) for realistic network sizes are not provided by us.
  5. 5.
    Fusion process using PBS In [16], they successfully produced GHZ states by performing Bell projection on photons from different Bell pairs using polarized beam splitting. This was done without considering the phase. Our framework requires strict phase control. This fusion with precise phase control needs to be empirically demonstrated.
  6. 6.
    Classical-quantum interface and operational integration. The transition between classical data and the quantum blockchain register is described conceptually, but a full operational specification (including classical-quantum interaction layers, transaction validation mechanisms, revocation models, and hybrid network protocols) is left for future work.